factor the trinomial
s^2-3sd-10d^2
the answer is?

Answer by jorel1380(3719)

s^2-3sd-10d^2 = (s-5d)(s+2d) = s^2-5ds+2ds-10d^2 = s^2-3ds-10d^2.